40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 21 18:45:50 2010, Koji, Update, PSL, Found the beat at 1064nm 
    Reply  Thu Oct 21 19:10:39 2010, Aidan, Update, PSL, Found the beat at 1064nm 
       Reply  Fri Oct 22 01:23:13 2010, Koji, Update, PSL, Found the beat at 1064nm 101021_beat.pdf
Message ID: 3759     Entry time: Fri Oct 22 01:23:13 2010     In reply to: 3756
Author: Koji 
Type: Update 
Category: PSL 
Subject: Found the beat at 1064nm 

[Koji / Suresh]

We worked on the 1064 beating a bit more.

- First of all, FSS and FSS SLOW servo were disabled not to have variating Xtal temp for the PSL.

- The PSL Xtal temp (T_PSL) was scanned from 22deg-45deg while we search the Xtal temp (T_Xend) for the Xend laser to have the beat freq well low (f<30MHz).
The pumping current for each laser was I_PSL = 2.101 [A] and I_Xend = 2.000 [A]

For a certain T_PSL, we found multiple T_Xend because the freq of the laser is not a monotonic function of the Xtal temperature. (see the innolight manual).

T_Xend to give us the beating was categorized in the three sets as shown in the figure. The set on "curve2" is the steadiest one. (Use this!)
The trends were quite linear but the slope was slightly off from the unity.

- T_PSL was scanned to see the trend of the PMC output.

The PMC was sometimes locked to the mode with lower transmission (V_PMCT ~ 3.0V).
When T_PSL ~ 31deg we consistently locked the PMC at higer transmission (V_PMCT ~ 5.3V).

At the moment we decided the operating point of T_PSL = 32.25 deg, V_PMCT = 5.34, where we found the beat at T_Xend=38.28deg.

- We cleaned up the PSL table more than how it was. Returned the tools to their original places.
The X-end laser was shut down and was left on the PSL table.

Kiwamu can move the X-end laser to the Xend and realign it.
Then we should be able to see the green beating quite easily.

Attachment 1: 101021_beat.pdf  71 kB  | Show | Show all
ELOG V3.1.3-